In a population of 100 individuals, we have 25 people that are homozygous AA, 50 people that are heterozygous Aa, and 25 people that are homozygous aa. If q is the frequency of a alleles in this population, calculate q. Write your answer as a decimal.
Answer #1

Answer: q = 0.5

Explanation:

AA25 = 25*2= 50 "A" alleles

AA50 = 50A & 50 a alleles

aa24 = 50 "a" alleles

Total alleles = 200

Total of "a" alleles = 50+50 = 100

Frequency of allele "a" = q = 100/200 = 0.5